# LINSCAN - A Linearity Based Clustering Algorithm

### v1

@@ Abstract
keywords:
Anisotropic DBSCAN , Linear clustering , OPTICS , Earthquakes , Faults
journal: Pattern Recognition Letters
[…]
[1]organization=Committee on Computational and Applied Mathematics, University of Chicago, city=Chicago, postcode=Postal Code, state=IL, country=USA
[…]
[2]organization=Institute of Geophysics and Planetary Physics, Scripps Institution of Oceanography, University of California San Diego, city=La Jolla, postcode=92093, state=CA, country=USA
[…]
[3]organization=Department of Structural Engineering, University of California San Diego, city=La Jolla, postcode=92093, state=CA, country=USA
[…]
[4]organization=Department of Mathematics and Halcioglu Data Science Institute, University of California San Diego, city=La Jolla, postcode=92093, state=CA, country=USA
@@ 1 Introduction
In contrast, DBSCAN and OPTICS iteratively generate clusters by leveraging a heuristic for the local behavior of clustered points. In particular, the designers equated clusters to connected regions of high density [1]. Thus, by identifying points whose local neighborhoods are highly dense, even with little prior knowledge about the local geometry of the data one can iteratively grow clusters from those points. The number of clusters then comes naturally from the geometry of the data itself, rather than being a parameter.
In this paper, we seek to leverage this characterization of clusters using a clustering metric other than Euclidean distance. In particular, we propose an algorithm that can distinguish between multiple quasi-linear clusters that may be closely spaced but have nearly orthogonal covariances. This is motivated in particular by the need to identify and map seismically active faults given a catalog of precisely located earthquakes, an important problem in geophysics [2, 3, 4]. In addition, the potential of the algorithm is not limited to geophysics, but it may also help identify the linear spatial patterns of other natural features such as soil and airborne pollution, and man-made directional patterns including roads and hiking trials [5, 6, 7].
@@ 1.1 Motivating Problem
We wish to isolate quasi-linear clusters (QLCs) in point clouds and distinguish clusters that are geometrically close, or possibly overlap, but have different orientations. This problem arises e.g. in geophysics when one attempts to identify active seismogenic faults based on epicentral locations of micro-earthquakes [8, 2, 4]. While faults are three-dimensional quasi-planar surfaces, in appropriate projections they appear as linear features, so that the associated locations of micro-earthquakes can be recognized as quasi-linear features after accounting for noise.
To highlight the deficiency of existing algorithms for this task, consider a synthetic data set shown in Figure 1(a). The data set includes QLCs, some of which intersect each other (e.g., see around coordinate (-.4,-.6)), as well as irregularly shaped clusters and ”background noise”. Figure 1(b) shows the results obtained by applying DBSCAN (described in Section 2) to the data. Note that the output includes both linear and irregular clusters, some QLCs are conjoined with irregular clusters, and QLCs that are geometrically close (e.g., intersecting or overlapping) are considered to be part of the same cluster.
Figure 1: Test Data
@@
(b) DBSCAN Results
1.2 Synopsis of this study
  1. 1.
    We have designed an algorithm that can be used to identify linear clusters in a point cloud without losing the stability guarantees of well-established clustering algorithms like DBSCAN and OPTICS.
  2. 2.
    We discuss possible ways to generalize DBSCAN and OPTICS to apply them to more specialized tasks. In particular, we discuss that as long as symmetry is maintained, non-metric distance functions can be used without losing the stability properties of DBSCAN, such as invariance of clustering behavior under permutations of the order of the points.
  3. 3.
    We compare our framework to ADCN [9], a previous attempt at applying DBSCAN to a similar task, and discuss how the design of ADCN leads to the shape and number of clusters being sensitive to changes in the order of the points. We then contrast this with LINSCAN, which is invariant to the ordering of the points for clustering.
  4. 4.
    We design a distance function related to the KL divergence and prove that while it is not a metric, it only fails to be one polynomially, meaning that clusters in this metric are spatially dense.
@@ 1.3 Notation
Here we summarize the notation that will be used throughout the rest of this paper:
  1. 1.
    For $\epsilon>0$, we let $B_{\epsilon}(x)$ be the open ball of radius $\epsilon$ centered at $x$ (in the standard Euclidean norm).
  2. 2.
    For finite $E\subseteq\mathbb{R}^{d}$,
    1. (a)
      $\mu_{E}\in\mathbb{R}^{d}$ is the sample mean of $E$.
    2. (b)
      $\Sigma^{2}_{E}\in\mathbb{R}^{d\times d}$ is the sample covariance matrix of $E$.
    Given $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ with $\Sigma$ symmetric positive definite, $\mathcal{N}(\mu,\Sigma)$ is the multivariate Gaussian distribution with mean $\mu$ and covariance $\Sigma$.
  3. 3.
    We let $\left\|A\right\|_{F}:=\mathrm{tr}(A^{T}A)$ denote the Frobenius norm.
  4. 4.
    For positive definite $A$, $\left\|x\right\|_{A}:=\sqrt{x^{T}Ax}$ is the elliptic norm defined by $A$.
  5. 5.
    For general matrix $A$ we let $|A|$ denote the determinant of $A$.
  6. 6.
    We let $\mathbb{P}(\mathcal{X})$ be the set of probability distributions over $\mathcal{X}$ and let $\mathcal{P}(\mathcal{X})$ be the power set of $\mathcal{X}$.
We begin by summarizing the most popular clustering algorithms, namely DBSCAN and OPTICS, to provide context for readers lacking a background in clustering theory. Those with a familiarity with DBSCAN and OPTICS can skip directly to section 3.
@@ 2.1 DBSCAN
In what follows assume $X=\left\{x_{1},...,x_{m}\right\}\subseteq\mathbb{R}^{d}$ is a point cloud and let $\epsilon>0$ and $\mathrm{minPts}\in\mathbb{N}$ be two parameters. We say $x\in X$ is a core point if #$\#(B_{\epsilon}(x)\cap X)>\mathrm{minPts},$ where $B_{\epsilon}(x)$ is the ball of radius $\epsilon$ about $x$.
@@
As a result, core reachability is an equivalence relation. DBSCAN then defines clusters to simply be equivalence classes under this relation, with clusters containing fewer than $\mathrm{minPts}$ points being labelled as noise. Algorithm 1 in the appendix provides a pseudocode description of how this is done.
@@ 2.2 OPTICS
OPTICS acts as a generalization of DBSCAN, improving its robustness on datasets with regions of various densities and partially abstracting away the $\epsilon$ parameter [10]. The most popular and effective implementation of OPTICS takes in three parameters: $\epsilon$, $\mathrm{minPts}$, and $\xi$, although $\epsilon$ is optional and only serves to shorten the run-time of the algorithm.
@@
For $p,o\in X$, we define the reachability distance from $o$ to $p$ as $d_{\mathrm{reach}}(p|o)=\max\left\{d_{\mathrm{core}}(o),\left\|p-o\right\|\right\}.$
The reachability distance describes the minimum $\epsilon$ such that $o$ is considered a core point and $p$ is contained in an $\epsilon$-neighborhood of $o$. Note that this can be infinite if $d_{\mathrm{core}}(o)=\infty$. OPTICS proceeds to develop a priority queue using a process described in the appendix in Algorithms 2 and 3.
@@ 2.3 Related Work
The choice to use Euclidean distance with DBSCAN/OPTICS is arbitrary. The stability of the algorithm only depends on the fact that the distance function is symmetric and non-negative. Importantly, the function does not need to satisfy the triangle inequality [e.g., 11, p. 8], which allows us to work with non-metrics.
The idea of extending DBSCAN/OPTICS to domains where we seek linearity is not entirely new. Previously, an algorithm called ADCN was developed to solve this problem by redefining the search neighborhoods from circles to ellipses whose eccentricity reflects the local covariance of the point [9]. In practice, ADCN performs as well as DBSCAN in many tasks and performs better in cases where clusters are locally linear in otherwise highly noisy datasets.
However, ADCN is not well-suited for our task in particular because it does not provide the desired separation of adjacent or intersecting QLCs. On the contrary, it can produce artifacts around the intersection areas, say for a T-shaped intersection as in Figure 2. Furthermore, the point selection process in ADCN is non-symmetric, meaning that in certain cases the clustering behavior may be unstable to permutations of the points. Figures 2(d) and 2(e) show two runs of ADCN on the same dataset with the same parameters, but with the dataset in a different order. Note how sensitive the behavior of the algorithm is to the order of the points. Our proposed algorithm performs more stably, as demonstrated below.
@@ 3.1 The Embedding and Distance
LINSCAN seeks to keep the advantages DBSCAN provides while applying it to the task of distinguishing QLCs. To do this, we embed data points into $\mathbb{P}(\mathbb{R}^{d})$, the space of probability measures on $\mathbb{R}^{d}$, and then cluster the data using a notion of distance between distributions. Letting $\mathrm{eccPts}\in\mathbb{N}$ and defining $R^{m}(x)$ be the $m$-nearest neighbors to $x$ in $X$, we define a mapping
@@
Thus, we embed each point in the dataset as the normal distribution best approximating its $\mathrm{eccPts}$-nearest neighbors, which allows us to cluster the points based on the local covariance of the data. Note how we rescale the covariance matrix to have maximal eigenvalue of 1.
@@
$D(P,Q)$ $=\frac{1}{2}\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|_{F}+\frac{1}{2}\left\|\Sigma_{P}^{-1/2}\Sigma_{Q}\Sigma_{P}^{-1/2}-I\right\|_{F}+\frac{1}{\sqrt{2}}\left\|\mu_{P}-\mu_{Q}\right\|_{\Sigma_{Q}^{-1}}+\frac{1}{\sqrt{2}}\left\|\mu_{P}-\mu_{Q}\right\|_{\Sigma_{P}^{-1}}$
where $P=\mathcal{N}(\mu_{P},\Sigma_{P})$ and $Q=\mathcal{N}(\mu_{Q},\Sigma_{Q})$ for positive definite $\Sigma_{P}$ and $\Sigma_{Q}$. Note that this function is symmetric and $D(P,Q)=0$ if and only if $P=Q$. Although $D$ does not satisfy the triangle inequality and is thus not a metric, later we will discuss an approximate form of the triangle inequality that $D$ does satisfy. Note that by choosing to normalize the covariances as above, we have
$D(P,Q)\geq\sqrt{2}\left\|\mu_{P}-\mu_{Q}\right\|_{2}$
Thus, points can be efficiently disqualified from consideration without having to calculate the more expensive matrix terms if the means are sufficiently far apart, which can be used to improve the run-time of the algorithm.
Once the points have been embedded as distributions, we run OPTICS on $\mathcal{P}=\left\{P_{i}\right\}_{i=1}^{m}$ with Euclidean distance replaced by $D(\cdot,\cdot)$, and cluster $X$ based on the results. The full process is described in Algorithm 4 (see appendix).
3.2 Motivating the Definition of $D$
@@
$KL(P|Q)$ $=\frac{1}{2}\log\frac{\left|\Sigma_{Q}\right|}{\left|\Sigma_{P}\right|}+\frac{1}{2}\mathrm{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)+\frac{1}{2}(\mu_{P}-\mu_{Q})^{T}\Sigma_{Q}^{-1}(\mu_{P}-\mu_{Q})$
One can show (see appendix A) that if $\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|_{F}<1$, then
$KL(P|Q)$ $=\frac{1}{4}\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|_{F}^{2}+o\left(\mathrm{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)+\frac{1}{2}(\mu_{P}-\mu_{Q})^{T}\Sigma_{Q}^{-1}(\mu_{P}-\mu_{Q})$
@@
We note that other metrics, in particular Wasserstein-2 distance, also have a closed form between Gaussians. While this is a metric, the distance between the means and covariances are independent, whereas $D$ incorporates the Mahalanobis distance and penalizes differences in mean more heavily in directions orthogonal to the local linearity of the point. Furthermore, the Wasserstein-2 distance penalizes differences in covariances at most polynomially in the magnitude of the eigenvalues, whereas $D$ penalizes orthogonal covariance inversely to the size of the minimum eigenvalues for high eccentricity clusters. This ensures that two points with slight deviations in covariance direction will be far apart in $D$, and thus these points will not fall into the same cluster.
3.3 Local Behavior of $D$
While $D$ does not satisfy the full triangle inequality, one can show that it satisfies a slightly relaxed version. We utilize the matrix commutator $[\cdot,\cdot]:\mathbb{R}^{d\times d}\times\mathbb{R}^{d\times d}\rightarrow\mathbb{R}^{d\times d}$, which measures the degree to which two matrices commute via $[A,B]:=AB-BA.$
Theorem 1

### v2

@@ 1 Introduction
In contrast, DBSCAN and OPTICS iteratively generate clusters by leveraging a heuristic for the local behavior of clustered points. In particular, the designers equated clusters to connected regions of high density (dbscan). Thus, by identifying points whose local neighborhoods are highly dense, even with little prior knowledge about the local geometry of the data, one can iteratively grow clusters from those points. The number of clusters then comes naturally from the geometry of the data itself, rather than being a parameter.
In this paper, we seek to leverage this characterization of clusters using a clustering metric other than Euclidean distance. In particular, we propose an algorithm that can distinguish between multiple quasi-linear clusters that may be closely spaced but have nearly orthogonal covariances. This is motivated in particular by the need to identify and map seismically active faults given a catalog of precisely located earthquakes, an important problem in geophysics (fi21; zou23; shelly2023fracture). In addition, the potential of the algorithm is not limited to geophysics, but it may also help identify the linear spatial patterns of other natural features such as soil and airborne pollution, and man-made directional patterns including roads and hiking trials (barden1963; isaaks1989; ADCN2).
An application of the algorithm described in this paper can be found in zou23, wherein it was used to identify slip faults at many scales to investigate the distribution of dihedral angles of conjugate faults in the Anza-Borrego Shear Zone.
@@ 1.1 Motivating Problem
We wish to isolate quasi-linear clusters (QLCs) in point clouds and distinguish clusters that are geometrically close, or possibly overlap, but have different orientations. Quasi linear clusters are a cluster of points where: 1) each point is within $\epsilon$ of some other point in the cluster, 2) the total cluster has a nearly singular covariance matrix. This problem arises, for example, in geophysics, when one attempts to identify active seismogenic faults based on epicentral locations of microearthquakes (cochran2020activation; fi21; shelly2023fracture). Although faults are three-dimensional quasi-planar surfaces, in appropriate projections they appear as linear features, so that the associated locations of micro-earthquakes can be recognized as quasi-linear features after accounting for noise.
As an example of this task, consider a synthetic data set shown in Figure 1. The data set includes QLCs, some of which intersect each other (e.g., see around coordinate (-.4,-.6)), as well as irregularly shaped clusters and "background noise." Note how LINSCAN is capable of separating even spatially dense clusters into their component QLCs, in contrast to DBSCAN. 1(c) shows the results after an optional post-processing step we discuss at the start of section 5.
Figure 1: Synthetic Data
@@
(b) LINSCAN Results
(c) Filtered Results
(d) DBSCAN Results
1.2 Contributions
  1. a.
    We design an algorithm that can be used to identify quasi-linear clusters in a point cloud without losing the stability guarantees of well-established clustering algorithms like DBSCAN and OPTICS.
  2. b.
    We compare our framework to ADCN (ADCN), a previous attempt at applying DBSCAN to a similar task, and discuss how the design of ADCN leads to the shape and number of clusters being sensitive to changes in the order of the points. This is in contrast to LINSCAN, which is invariant to the ordering of the points for clustering.
  3. c.
    We prove that while our distance measure is not a metric, it satisfies positivity and symmetry on the space of Gaussian distributions (see Lemma 3.1), and a slightly relaxed form of the triangle inequality (see Theorem 3.2). These results combine to mean that clusters in this metric are stable under the order of the points and are spatially dense.
@@ 1.3 Notation
Here we summarize the notation that will be used throughout the rest of this paper. For finite $E\subseteq{\mathbb{R}}^{d}$, we let $\mu_{E}$ and $\Sigma_{E}$ be the sample mean and covariance of $E$. For positive definite $A$, ${\left\|x\right\|}_{A}:=\sqrt{x^{T}Ax}$ is the elliptic norm defined by $A$. Finally, a QLC is a set $S$ satisfying:
  1. a.
    $\forall x\in S$, $\exists y\in S\setminus\left\{x\right\}$ such that $\|x-y\|<\epsilon$ for some small $\epsilon$,
  2. b.
    the covariance $\Sigma_{S}$ satisfies $\tau<\mathrm{cond}_{2}(\Sigma_{S}):=\frac{\sigma_{\mathrm{max}}(\Sigma_{S})}{\sigma_{\mathrm{min}}(\Sigma_{S})}$ for some large $\tau$.
@@ 2.1 DBSCAN
In what follows, assume $X=\left\{x_{1},...,x_{m}\right\}\subseteq{\mathbb{R}}^{d}$ is a point cloud and let $\epsilon>0$ and $\mathrm{minPts}\in{\mathbb{N}}$ be two parameters. We say $x\in X$ is a core point if $(B_{\epsilon}(x)\cap X)>\mathrm{minPts},$ where $B_{\epsilon}(x)$ is the ball of radius $\epsilon$ about $x$.
@@
As a result, core reachability is an equivalence relation. DBSCAN then defines clusters to simply be equivalence classes under this relation, with clusters containing fewer than $\mathrm{minPts}$ points being labeled as noise. Algorithm 1 in the supplemental documents provides a pseudocode description of how this is done.
@@ 2.2 OPTICS
OPTICS acts as a generalization of DBSCAN, improving its robustness on datasets with regions of various densities and partially abstracting away the $\epsilon$ parameter (OPTICS). The most popular and effective implementation of OPTICS takes in three parameters: $\epsilon$, $\mathrm{minPts}$, and $\xi$, although $\epsilon$ is optional and only serves to shorten the run-time of the algorithm.
@@
For $p,o\in X$, we define the reachability distance from $o$ to $p$ as $d_{\mathrm{reach}}(p|o)=\max\left\{d_{\mathrm{core}}(o),{\left\|p-o\right\|}\right\}.$ The reachability distance describes the minimum $\epsilon$ such that $o$ is considered a core point and $p$ is contained in an $\epsilon$-neighborhood of $o$. Note that this can be infinite if $d_{\mathrm{core}}(o)=\infty$. OPTICS proceeds to develop a priority queue using a process described in the supplemental document in Algorithms 2 and 3.
@@ 2.3 Related Work
The choice to use Euclidean distance with DBSCAN/OPTICS is arbitrary. The stability of the algorithm only depends on the fact that the distance function is symmetric and non-negative. Importantly, the function does not need to satisfy the triangle inequality (e.g., khamsi2011introduction, p. 8), which allows us to work with non-metrics.
Anisotropic DBSCAN: The idea of extending DBSCAN/OPTICS to domains where we seek linearity is not entirely new. Previously, an algorithm called ADCN was developed to solve this problem by redefining the search neighborhoods from circles to ellipses whose eccentricity reflects the local covariance of the point (ADCN). In practice, ADCN performs as well as DBSCAN in many tasks and performs better in cases where clusters are locally linear in otherwise highly noisy datasets.
However, ADCN is not well-suited for our task in particular because it does not provide the desired separation of adjacent or intersecting QLCs. On the contrary, it can produce artifacts around the intersection areas, say for a T-shaped intersection as in Figure 3. Furthermore, the point selection process in ADCN is non-symmetric, meaning that in certain cases the clustering behavior may be unstable to permutations of the points. Figures 3(d) and 3(e) show two runs of ADCN on the same dataset with the same parameters but with the dataset in a different order. Note how sensitive the behavior of the algorithm is to the order of the points. Our proposed algorithm performs more stably, as demonstrated below.
Anisotropic Kernels and Spectral Clustering: There are a large number of kernel method algorithms that use anisotropic kernels and local Mahalanobis distances to define similarity, see for example wang2007weighted; talmon2013empirical; arias2017spectral; lahav2019mahalanobis; cheng2020two; peterfreund2020local. In practice, these can capture a similar notion of local similarity to our proposed approach and have been used for spectral clustering. For example, arias2017spectral considers a similar problem to ours in clustering data that arises from intersecting manifolds.
However, regardless of the kernel similarity, spectral clustering and k-means (or another clustering algorithm) in the latent space fail in our noisy setting, where most of the points do not belong to any cluster. This is because k-means and spectral clustering algorithms perform poorly for data sets that are not a union of well-separated clusters (either in the original space or feature space of the kernel) little2020path, which was the motivation for the initial development of DBSCAN. There exist DBSCAN-like spectral clustering algorithms that are robust to outliers by using path-based similarity chang2008robust; little2020path, but these algorithms have no bias towards QLCs or other degenerate clusters. For these reasons, we do not include explicit comparions to these methods in this manuscript and restrict ourselves to DBSCAN/OPTICS based algorithms.
@@ 3.1 The Embedding and Distance
LINSCAN seeks to keep the advantages of DBSCAN while being applicable to the task of distinguishing QLCs. To do this, we embed data points into $\mathbb{P}({\mathbb{R}}^{d})$, the space of probability measures on ${\mathbb{R}}^{d}$, and then cluster the data using a notion of distance between distributions.
LINSCAN has 3 required parameters $\mathrm{minPts}$, $\mathrm{eccPts}$, and $\xi$ and one optional parameter $\epsilon$. $\mathrm{minPts}$, $\xi$, and $\epsilon$ are identical to the corresponding parameters in OPTICS, but $\mathrm{eccPts}$ is a parameter specific to LINSCAN which determines how we form the distributions we use for clustering. Letting $R^{m}(x)$ be the $m$-nearest neighbors to $x$ in $X$, we define a mapping
@@
Thus, we embed each point in the dataset as the normal distribution best approximating its $\mathrm{eccPts}$-nearest neighbors, which allows us to cluster the points based on the local covariance of the data. Note that we rescale the covariance matrix to have maximal eigenvalue of 1.
@@
$\quad\quad+\frac{1}{\sqrt{2}}{\left\|\mu_{P}-\mu_{Q}\right\|}_{\Sigma_{Q}^{-1}}+\frac{1}{\sqrt{2}}{\left\|\mu_{P}-\mu_{Q}\right\|}_{\Sigma_{P}^{-1}}$
where $P=\mathcal{N}(\mu_{P},\Sigma_{P})$ and $Q=\mathcal{N}(\mu_{Q},\Sigma_{Q})$ for positive definite $\Sigma_{P}$ and $\Sigma_{Q}$. Note that this function is symmetric and $D(P,Q)=0$ if and only if $P=Q$. Although $D$ does not satisfy the triangle inequality and is thus not a metric, later we will discuss an approximate form of the triangle inequality that $D$ does satisfy (see Theorem 3.2). Note that by choosing to normalize the covariances as above, we have
$D(P,Q)\geq\sqrt{2}{\left\|\mu_{P}-\mu_{Q}\right\|}_{2}$ (1)
Thus, points can be efficiently disqualified from consideration without having to calculate the more expensive matrix terms if the means are sufficiently far apart, which can be used to improve the run-time of the algorithm. This is, in particular, how we utilize $\epsilon$, as this means we can filter out pairs points using standard spatial methods (KD-Trees, etc.) in Euclidean space to filter out points that are sufficiently far apart without having to compute their distance in our distance measure.
Once the points have been embedded as distributions, we run OPTICS on ${\mathcal{P}}=\left\{P_{i}\right\}_{i=1}^{m}$ with Euclidean distance replaced by $D(\cdot,\cdot)$, and cluster $X$ based on the results. The full process is described in Algorithm 4 (see supplemental document).
3.2 Motivation of Distance Measure
@@
$KL(P|Q)$ $=\frac{1}{2}\log\frac{\left|\Sigma_{Q}\right|}{\left|\Sigma_{P}\right|}+\frac{1}{2}\operatorname{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)+\frac{1}{2}(\mu_{P}-\mu_{Q})^{T}\Sigma_{Q}^{-1}(\mu_{P}-\mu_{Q})$
One can show (see supplemental document) that if ${\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|}_{F}<1$, then
$KL(P|Q)$ $=\frac{1}{4}{\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|}_{F}^{2}+\frac{1}{2}(\mu_{P}-\mu_{Q})^{T}\Sigma_{Q}^{-1}(\mu_{P}-\mu_{Q})$
$\quad\quad+o\left(\operatorname{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
@@
We note that other metrics, in particular Wasserstein-2 distance, also have a closed form between Gaussians. While this is a metric, the distance between the means and covariances are independent, whereas $D$ incorporates the Mahalanobis distance and penalizes differences in mean more heavily in directions orthogonal to the local linearity of the point. Furthermore, the Wasserstein-2 distance scales polynomial in the magnitude of the eigenvalues of the covariance matrices as the angles diverge, whereas $D$ penalizes orthogonal covariance inversely to the size of the minimum eigenvalues for high eccentricity clusters. This ensures that two points with large deviations in covariance direction will be far apart in $D$, even if spatially close to one another, and thus these points will not fall into the same cluster.
3.3 Behavior of Distance Measure
Our distance measure is not a metric. However, in the case of Gaussians, it satisfies the properties of symmetry and separation of points in general, and, as we will show in Theorem 3.2, it satisfies a relaxed form of the triangle inequality.
Lemma 3.1.
Let $P={\mathcal{N}}(\mu_{P},\Sigma_{P})$ and $Q={\mathcal{N}}(\mu_{Q},\Sigma_{Q})$ be Gaussians. Then,
  1. a.
    $D$ is symmetric, meaning
    $D(P,Q)=D(Q,P)$
  2. b.
    $D(P,Q)=0$ iff $P=Q$ (in particular $D(P,P)=0$)
The proof of this lemma is in the appendix. While $D$ does not satisfy the full triangle inequality, one can show that it satisfies a slightly relaxed version. We utilize the matrix commutator $\cdot,\cdot]:\mathbb{R}^{d\times d}\times\mathbb{R}^{d\times d}\rightarrow\mathbb{R}^{d\times d}$, which measures the degree to which two matrices commute via [A,B]:=ABBA.[A,B]:=AB-BA.
Theorem 3.2.